# algorithmic

### v1

@@ 2 Related Work
DR methods (Dudik et al., 2011; 2014; Wang et al., 2017; Farajtabar et al., 2018) strike a balance between the biased, low-variance DM and the unbiased, high-variance IPS by incorporating both. It can also be interpreted as using control variates within an IPS framework (Thomas & Brunskill, 2016; Veness et al., 2011). If either IPS or DM is unbiased, DR methods are guaranteed to be unbiased (Dudik et al., 2011). Much of the follow-up work on DR methods has focused on mitigating the detrimental effects of variance or extreme probabilities from the IPS component. For example, SnDR utilizes SnIPS (Swaminathan & Joachims, 2015b), DR-SWITCH truncates IPS using a threshold (Wang et al., 2017), and DR-Shrinkage finds a weighting strategy by optimizing a sharp bound on mean squared error (Su et al., 2019a). While much of the prior work has emphasized improving the IPS component of DR estimators and their asymptotical guarantees, our research targets enhancing the robustness of conditional reward estimators, especially under large shifts and limited data. Our key insight is to view this problem as a covariate shift and utilize robust regression to train the conditional reward estimator. Our proposed method can also be naturally incorporated into existing methods like DMs and DRs by offering a distributionally robust reward estimator.
@@
Distributionally robust optimization (DRO) (Ben-Tal et al., 2009; Słowik & Bottou, 2022), a prevalent approach to handling distribution shifts, introduces the idea of optimizing with respect to the most adverse scenarios over a defined family of distributions. This has led to different variants such as Wasserstein DRO (Blanchet et al., 2021). A few studies have explored distributional robustness in policy evaluation. Si et al. (2020) and Kallus et al. (2022) augmented evaluation robustness by conservatively estimating reward mapping with KL-divergence uncertainty sets. However, in these methods, the uncertainty set covers all possible shifts in the joint distribution of context, action, and reward, which does not characterize the covariate shift specifically. As a result, the estimator can be overly robust and less effective. Robust regression (Chen et al., 2016; Liu & Ziebart, 2017) tailored distributional robustness methods specifically for the setting of covariate shift. It constructs a predictor that approximately matches training data statistics but is otherwise the most uncertain on the testing distribution, by minimizing the worst-case expected target loss and obtaining a parametric form of the predicted output labels’ probability distributions. We are the first to employ robust regression methods for policy evaluation, enabling the refinement of direct methods and the further enhancement of doubly robust estimators.
@@ 3 Problem Description
$V^{T}=\mathbb{E}_{\bm{x}\sim P_{t}(\bm{x}),a\sim\pi(a|\bm{x})}\left[r_{\bm{x},a}\right],$ (1)
@@
Figure 1: Data generating process for policy evaluation under general covariate shift. $X$, $A$, and $R$ represent context, action, and reward variables, respectively. Logging data distribution $P_{s}(x)\beta(a|\bm{x})$ is marked with solid purple lines, and target data distribution $P_{t}(x)\pi(a|\bm{x})$ is marked with dashed red lines. Black lines represent that reward is generated condition on given pairs of contexts and actions.
@@ 4.1 The Distributionally Robust Learning Formulation
Derived Predictive Form. If we also choose the base distribution to be Gaussian as $f_{0}\sim\mathcal{N}(\mu_{0},\sigma^{2}_{0})$, we have the following form for the predictor $\hat{f}$: $\hat{f}_{\bm{\theta}}(\bm{x},a)\sim\mathcal{N}(\mu_{\bm{\theta}}(\bm{x},a),\sigma_{\bm{\theta}}^{2}(\bm{x},a))$ (The full derivation can be found in Appendix B):
@@ 4.2 The Proposed Algorithm for Parameter Learning
We summarize the proposed robust regression approach for reward prediction in Figure 2 (a). After deriving the predictive form, we here present how to learn model parameters $\bm{\theta}$. With the predictive form of $\hat{f}$ in Equation 5, we can plug it back into the Lagrangian form of the constrained optimization problem in Equation 3 and Equation 4, and obtain the target loss to be equivalent as maximizing the log-likelihood over the target distribution: $\mathcal{L}(\bm{\theta})=\mathbb{E}_{P_{t}(\bm{x},a,r)}\left[\log\hat{f}_{\bm{\theta}}(\bm{x},a)\right]$. Given the form of $\hat{f}$, this objective is convex w.r.t $\bm{\theta}$. By taking the derivative of $\mathcal{L}(\bm{\theta})$ w.r.t the $\theta_{r}$ and $\bm{\theta}_{\bm{x}}$, we can obtain the following gradient form, where $S_{\text{batch}}$ is the set for data mini-batch. Standard gradient-based learning can then be applied for learning the parameters. The training procedure is summarized in Algorithm 1. Detailed derivation can be found in Appendix B.
@@ 5 Methods
In this section, we first provide an overview of how our robust regression approach can be integrated into existing policy evaluation methods as a reward model. A detailed comparison between classical OPE and the setting of policy evaluation under general covariate shift considered in this paper is illustrated in Figure 2 (b). We then establish upper bounds for the bias of these methods. We will refer to our methods using the notation PS and GCS as suffixes, representing policy shift and general covariate shift, respectively. Specifically, the proposed methods DM-PS, DR-PS, DM-GCS, DR-GCS, SnDR-PS and SnDR-GCS are formulated as follows.
@@
$\hat{V}_{\text{DM-PS}}=\frac{1}{|S|}\sum_{\bm{x}\in S}\mathbb{E}_{a\sim\pi}\left[\mu^{\text{PS}}_{\bm{\theta}}(\bm{x},a)\right].$ (8)
@@
DM-GCS. Under general covariate shift, we set $\mathcal{W}(\bm{x},a)$ to be $\frac{P_{s}(\bm{x},a)}{P_{t}(\bm{x},a)}$ in robust regression and plug it into the standard DM to get “Direct Method with General Covariate Shift” (DM-GCS) estimator. We also plug $\mathcal{W}(\bm{x},a)=\frac{P_{s}(\bm{x},a)}{P_{t}(\bm{x},a)}$ into IPS to implement IPS-GCS as a baseline. More details can be found in Appendix A.
$\hat{V}_{\text{DM-GCS}}=\frac{1}{|S|}\sum_{\bm{x}\in S}\mathbb{E}_{a\sim\pi}\left[\mu^{\text{GCS}}_{\bm{\theta}}(\bm{x},a)\right],$ (10)
@@
$\hat{V}_{\text{DR-GCS}}=\frac{1}{|S|}\sum_{(\bm{x},a,r)\in S}\left[\frac{(r-\mu_{\bm{\theta}}^{\text{GCS}}(\bm{x},a))P_{t}(\bm{x},a)}{P_{s}(\bm{x},a)}\right]+\hat{V}_{\text{DM-GCS}}.$ (11)
@@ 5.1.1 Bias Analysis
where $\mathcal{F}$ is the function class of $f$ with $\sup_{f,f^{\prime}\in\mathcal{F},\bm{x},a}|f(\bm{x},a)-f^{\prime}(\bm{x},a)|\leq M$ and a Rademacher complexity of $-^\hat{\mathfrak{R}}$, and $\eta_{1}$ upper bounds the gradient of optimization algorithm when it converges, as shown in Equation 31.
@@
$\mathbb{E}_{P_{t}(\bm{x},a,r)}\left[(r-\hat{\mu}(\bm{x},a))\right]\leq\left[mC\left[\frac{1}{n}\sum_{i=1}^{n}\sigma^{2}(\bm{x}_{i},a_{i})+\eta_{1}+4M\hat{\mathfrak{R}}(\mathcal{F})+3M^{2}\sqrt{\frac{\log\frac{2}{\delta}}{2n}}\right]+(1-m)H\right]^{1/2}:=\epsilon.$ (14)
@@
$\left|\mathbb{E}_{P_{t}(\bm{x},a,r)}\left[\hat{V}_{\mathrm{DR-PS/DR-GCS}}-V^{T}\right]\right|\leq\frac{C\eta_{1}}{l}+2CM\hat{\mathfrak{R}}(\mathcal{F})+3CM\sqrt{\frac{\log\frac{2}{\delta}}{2n}}+\epsilon.$ (15)
The proofs of Theorem 1 and 2 can be found in Appendix C.
Robustness under Large Shift. The above bias analysis shows that the proposed methods remain robust under large distribution shifts. Our bias analysis results consist of two parts, the first term is based on data with a non-zero density ratio where we reweight the target distribution into source distribution and use standard Rademacher complexity. The second term refers to the bounded error of the base distribution for target data that do not share support with the source data. In this case, the $m$ value, which is the percentage of target data points well-supported by the source distribution, represents the magnitude of the distribution shift. When the shift becomes larger, more proportion of the data will have a density ratio close to 0, and $m$ becomes smaller, corresponding to cases where the prediction relies more heavily on the base predictor $f_{0}$. Note that traditional ERM methods can have an arbitrarily bad performance when there is a severe non-overlapping issue between the source and target data distribution, such as large part of the target distribution is uncovered by the source distribution. Reweighting-based methods also become ineffective as the importance weights will have a zero denominator in a subset of data points. Our methods can effectively alleviate the issue as the bias of the proposed methods, which is shown in Equation 14 and Equation 15, will lean towards utilizing the base predictor $f_{0}$ and have a bounded error as $m$ becomes closer to 00. This makes our method more favorable than other methods under large shift settings, as further illustrated in the experiments section. In particular, we demonstrate how our method performs compared with other methods under different scales of shifting in Section 3. The experimental results show that our method remains robust and outperforms other methods, especially when the shift becomes large, as shown in Figure 4. Besides, we also provide a consistency analysis result for the settings where the density ratio is bounded away from 00 in Appendix C.1.
@@ 6.1 Experiment Settings
Datasets. In line with the experimental settings employed in previous studies (Dudik et al., 2014; Wang et al., 2017; Farajtabar et al., 2018; Su et al., 2019b; a), we conduct experiments on 9 UCI datasets by transforming the classification problems to the contextual bandits setting. Specifically, the context is set to be the input feature of the dataset, while the policy is a classification model and the action is the predicted class from the model. The reward is 1 if the predicted class matches the ground truth label and 0 otherwise. More details about how the dataset is created can be found in Appendix D.1, and the statistics of these datasets are summarized in Table 1 as follows.
@@
Covariate Shift on the Context. In our study, we examine two different types of context shifts within the data. We uniformly sample context for the target data, while manipulating the distribution of the context in the logging data in the following two ways to create covariate shift. 1) Gaussian covariate shift: We first apply Principal Component Analysis (PCA) to reduce the dimension. We then define a Gaussian distribution on the first principal component as the new data distribution. 2) Tweak-1 covariate shift ($\omega$): Following the logic of the Tweak-1 ($\rho$) logging policy, we assign a specific class of data a higher weight $\omega$ and all other classes with weight 1, and then do weighted sampling. Detailed information about the policies and data generation process can be found in Appendix D.1.
@@
Baselines. For our experimental comparisons, we consider several baseline methods, including Direct Method (DM), Inverse Propensity Score (IPS), Doubly Robust (DR) (Robins & Rotnitzky, 1995; Bang & Robins, 2005; Dudik et al., 2011), SnIPS and Self-Normalized Doubly Robust (SnDR) (Swaminathan & Joachims, 2015b). Also, we include a variant of our proposed methods DM (R), shown in Equation 23, as a baseline by setting the density ratio in the robust regression to 1. More details about the baselines, including SnIPS, SnDR, IPS-GCS, and DM (R), can be found in Appendix A.

### v2

@@ 2 Related Work
Off-Policy Evaluation for Contextual Bandits. Classic methods for OPE include direct methods (DM), inverse propensity scoring (IPS), and doubly robust (DR) methods that combine DM and IPS. DM can suffer from significant bias Dudik et al. (2011). IPS, though unbiased, exhibits high variance. DR methods (Dudik et al., 2011; 2014; Wang et al., 2017; Farajtabar et al., 2018) balance the strengths of both. If either IPS or DM is unbiased, DR methods are guaranteed to be unbiased (Dudik et al., 2011). Much of the follow-up work on DR methods has focused on mitigating the detrimental effects of variance or extreme probabilities from the IPS component. For example, SnDR utilizes SnIPS (Swaminathan & Joachims, 2015b), SWITCH estimator truncates IPS using a threshold (Wang et al., 2017), and DR with Optimistic Shrinkage (DRoS) finds a weighting strategy by optimizing a sharp bound on mean squared error (Su et al., 2019a). While much of the prior work has emphasized improving the IPS component of DR estimators and their asymptotical guarantees, our research targets enhancing the robustness of conditional reward estimators, especially under large shifts and limited data. Our key insight is to view this problem as a covariate shift problem and utilize robust regression to train the conditional reward estimator.
Covariate Shift and Distributional Robustness. Covariate shift has been tackled through methods such as importance weighting (Shimodaira, 2000; Sugiyama et al., 2007) and kernel mean matching (Gretton et al., 2009) in machine learning. In policy evaluation area, Uehara et al. (2020) introduced a DR-based estimator under covariate shift that uses a density ratio estimator between the source and target distributions. Their method focuses on the propensity score weighting formulation, and the IPS component in the estimator introduces high variance when extreme weights exist. In contrast, we tackle the problem orthogonally, utilizing a distributionally robust direct method that can be effortlessly combined with their method.
@@
There are several studies exploring distributionally robust optimization (DRO) (Ben-Tal et al., 2009; Słowik & Bottou, 2022; Blanchet et al., 2021) in policy evaluation. Si et al. (2020) and Kallus et al. (2022) augmented evaluation robustness by conservatively estimating reward mapping with KL-divergence uncertainty sets. However, in these methods, the uncertainty set covers all possible shifts in the joint distribution of context, action, and reward, which does not characterize the covariate shift specifically. As a result, the estimator can be overly robust and less effective. Another line of work utilize DRO for estimating confidence interval in the off-policy setting (Karampatziakis et al., 2020; Dai et al., 2020; Faury et al., 2020), following a generalized empirical likelihood approach (Duchi et al., 2021). The differences from our approach lie in the formulation of the DRO problem. Specifically, they do not use the conditional distribution of the target variable (the reward in our case) as the adversarial player, and the uncertainty set constraint is formulated by the perturbation of empirical distribution of the data with distance measured using f-divergence, rather than using feature matching as in our case. Robust regression (Chen et al., 2016; Liu & Ziebart, 2017) tailored distributional robustness methods specifically for the setting of covariate shift. It constructs a predictor that approximately matches training data statistics but is otherwise the most uncertain on the testing distribution by minimizing the worst-case expected target loss and obtaining a parametric form of the predicted output labels’ probability distributions. We are the first to employ robust regression methods for policy evaluation, enabling the refinement of direct methods and the further enhancement of doubly robust estimators.
Detailed Related Work. We provide a more detailed review of related work in Appendix A.
@@ 3 Problem Description
$\textstyle V^{T}=\mathbb{E}_{\bm{x}\sim P_{t}(\bm{x}),a\sim\pi(a|\bm{x})}\left[r_{\bm{x},a}\right],$ (1)
@@
Figure 1: Data generating process for policy evaluation under general covariate shift. $X$, $A$, and $R$ represent context, action, and reward variables, respectively. Logging data distribution $P_{s}(x)\beta(a|\bm{x})$ is marked with solid purple lines, and target data distribution $P_{t}(x)\pi(a|\bm{x})$ is marked with dashed red lines. Black lines indicate reward is generated by both contexts and actions.
@@ 4.1 The Distributionally Robust Learning Formulation
Derived Predictive Form. If we also choose the base distribution to be Gaussian as $f_{0}\sim\mathcal{N}(\mu_{0},\sigma^{2}_{0})$, we have the following form for the predictor $\hat{f}$: $\hat{f}_{\bm{\theta}}(\bm{x},a)\sim\mathcal{N}(\mu_{\bm{\theta}}(\bm{x},a),\sigma_{\bm{\theta}}^{2}(\bm{x},a))$ (The full derivation can be found in Appendix C):
@@
The Role of Gaussian Predictive Form. We assume a Gaussian base distribution and a quadratic feature constraint, which results in a Gaussian conditional target distribution $p(r|\bm{x},a)$. While this Gaussian assumption is generally valid for real-world examples and widely applied in ML literature, we can construct synthetic examples to violate such assumption. For instance, we can define the conditional target distribution as an exponential family with cubic terms over the exponential, corresponding to a cubic function feature constraint. In such cases, the standard bias and variance tradeoff analysis applies: using a model with Gaussian assumption in this situation leads to an overly simplified model, resulting in high bias and low variance. However, in practice, when little is known about the underlying data generation process, the robust regression model under Gaussian assumption is preferred due to its nice implementation properties and broad applicability. Further discussions on the Gaussian assumption is provided in Appendix C.1.
@@ 4.2 The Proposed Algorithm for Parameter Learning
We summarize the proposed robust regression approach for reward prediction in Figure 2 (a). After deriving the predictive form, we here present how to learn model parameters $\bm{\theta}$. With the predictive form of $\hat{f}$ in Equation 5, we can plug it back into the Lagrangian form of the constrained optimization problem in Equation 3 and Equation 4, and obtain the target loss to be equivalent as maximizing the log-likelihood over the target distribution: $\mathcal{L}(\bm{\theta})=\mathbb{E}_{P_{t}(\bm{x},a,r)}\left[\log\hat{f}_{\bm{\theta}}(\bm{x},a)\right]$. Given the form of $\hat{f}$, this objective is convex w.r.t $\bm{\theta}$. By taking the derivative of $\mathcal{L}(\bm{\theta})$ w.r.t the $\theta_{r}$ and $\bm{\theta}_{\bm{x}}$, we can obtain the following gradient form, where $S_{\text{batch}}$ is the set for data mini-batch. Standard gradient-based learning can then be applied for learning the parameters. The training procedure is summarized in Algorithm 1. Detailed derivation are in Appendix C.
@@ 5 Methods
In this section, we first provide an overview of how our robust regression approach can be integrated into existing policy evaluation methods as a reward model. A detailed comparison between classical OPE and the setting of policy evaluation under general covariate shift considered in this paper is illustrated in Figure 2 (b). We then establish upper bounds for the bias of these methods. We will refer to our methods using the notation PS and GCS as suffixes, representing policy shift and general covariate shift, respectively. Specifically, the proposed methods DM-PS, DR-PS, DM-GCS, DR-GCS, SnDR-PS and SnDR-GCS are formulated as follows. We also investigate the integration of our robust regression approach into other OPE methods including SWITCH (Wang et al., 2017) and DRoS (Su et al., 2020) in Appendix E.4.
@@
$\textstyle\hat{V}_{\text{DM-PS}}=\frac{1}{|S|}\sum_{\bm{x}\in S}\mathbb{E}_{a\sim\pi}\left[\mu^{\text{PS}}_{\bm{\theta}}(\bm{x},a)\right].$ (8)
@@
DM-GCS. Under general covariate shift, we set $\mathcal{W}(\bm{x},a)$ to be $\frac{P_{s}(\bm{x},a)}{P_{t}(\bm{x},a)}$ in robust regression and plug it into the standard DM to get “Direct Method with General Covariate Shift” (DM-GCS) estimator. We also plug $\mathcal{W}(\bm{x},a)=\frac{P_{s}(\bm{x},a)}{P_{t}(\bm{x},a)}$ into IPS to implement IPS-GCS as a baseline. More details can be found in Appendix B.
$\textstyle\hat{V}_{\text{DM-GCS}}=\frac{1}{|S|}\sum_{\bm{x}\in S}\mathbb{E}_{a\sim\pi}\left[\mu^{\text{GCS}}_{\bm{\theta}}(\bm{x},a)\right],$ (10)
@@
$\textstyle\hat{V}_{\text{DR-GCS}}=\frac{1}{|S|}\sum_{(\bm{x},a,r)\in S}\left[\frac{(r-\mu_{\bm{\theta}}^{\text{GCS}}(\bm{x},a))P_{t}(\bm{x},a)}{P_{s}(\bm{x},a)}\right]+\hat{V}_{\text{DM-GCS}}.$ (11)
@@ 5.1.1 Bias Analysis
where $\mathcal{F}$ is the function class of $f$ with $\sup_{f,f^{\prime}\in\mathcal{F},\bm{x},a}|f(\bm{x},a)-f^{\prime}(\bm{x},a)|\leq M$ and a Rademacher complexity of $\hat{\mathfrak{R}}$, and $\eta_{1}$ upper bounds the gradient of optimization algorithm when it converges, as shown in Equation 31.
@@
$\textstyle\mathbb{E}_{P_{t}(\bm{x},a,r)}\left[(r-\hat{\mu}(\bm{x},a))\right]\leq\left[mC\left[\frac{1}{n}\sum_{i=1}^{n}\sigma^{2}(\bm{x}_{i},a_{i})+\eta_{1}+4M\hat{\mathfrak{R}}(\mathcal{F})+3M^{2}\sqrt{\frac{\log\frac{2}{\delta}}{2n}}\right]+(1-m)H\right]^{1/2}:=\epsilon.$ (14)
@@
$\textstyle\left|\mathbb{E}_{P_{t}(\bm{x},a,r)}\left[\hat{V}_{\mathrm{DR-PS/DR-GCS}}-V^{T}\right]\right|\leq\frac{C\eta_{1}}{l}+2CM\hat{\mathfrak{R}}(\mathcal{F})+3CM\sqrt{\frac{\log\frac{2}{\delta}}{2n}}+\epsilon.$ (15)
The proofs of Theorem 1 and 2 can be found in Appendix D.
Robustness under Large Shift. The above bias analysis shows that the proposed methods remain robust under large distribution shifts. Our bias analysis results consist of two parts, the first term is based on data with a non-zero density ratio where we reweight the target distribution into source distribution and use standard Rademacher complexity. The second term refers to the bounded error of the base distribution for target data that do not share support with the source data. In this case, the $m$ value, which is the percentage of target data points well-supported by the source distribution, represents the magnitude of the distribution shift. When the shift becomes larger, more proportion of the data will have a density ratio close to 0, and $m$ becomes smaller, corresponding to cases where the prediction relies more heavily on the base predictor $f_{0}$. Note that traditional ERM methods can have an arbitrarily bad performance when there is a severe non-overlapping issue between the source and target data distribution, such as large part of the target distribution is uncovered by the source distribution. Reweighting-based methods also become ineffective as the importance weights will have a zero denominator in a subset of data points. Our methods can effectively alleviate the issue as the bias of the proposed methods, which is shown in Equation 14 and Equation 15, will lean towards utilizing the base predictor $f_{0}$ and have a bounded error as $m$ becomes closer to 00. This makes our method more favorable than other methods under large shift settings, as further illustrated in the experiments section. In particular, we demonstrate how our method performs compared with other methods under different scales of shifting in Section 3. The experimental results show that our method remains robust and outperforms other methods, especially when the shift becomes large, as shown in Figure 4. Besides, we also provide a consistency analysis result for the settings where the density ratio is bounded away from 00 in Appendix D.1.
@@ 6.1 Experiment Settings
Datasets. In line with the experimental settings employed in previous studies (Dudik et al., 2014; Wang et al., 2017; Farajtabar et al., 2018; Su et al., 2019b; a), we conduct experiments***The code for the experiments is available at https://github.com/guoyihonggyh/Distributionally-Robust-Policy-Evaluation-under-General-Covariate-Shift-in-Contextual-Bandits. on 9 UCI datasets by transforming the classification problems to the contextual bandits setting. Specifically, the context is set to be the input feature of the dataset, while the policy is a classification model and the action is the predicted class from the model. The reward is 1 if the predicted class matches the ground truth label and 0 otherwise. More details about how the dataset is created can be found in Appendix E.1, and the statistics of these datasets are summarized in Table 1 as follows.
@@
Covariate Shift on the Context. In our study, we examine two different types of context shifts within the data. We uniformly sample context for the target data, while manipulating the distribution of the context in the logging data in the following two ways to create covariate shift. 1) Gaussian covariate shift: We first apply Principal Component Analysis (PCA) to reduce the dimension. We then define a Gaussian distribution on the first principal component as the new data distribution. 2) Tweak-1 covariate shift ($\omega$): Following the logic of the Tweak-1 ($\rho$) logging policy, we assign a specific class of data a higher weight $\omega$ and all other classes with weight 1, and then do weighted sampling. Detailed information about the policies and data generation process can be found in Appendix E.1.
@@
Baselines. For our experimental comparisons, we consider several baseline methods, including Direct Method (DM), Inverse Propensity Score (IPS), Doubly Robust (DR) (Robins & Rotnitzky, 1995; Bang & Robins, 2005; Dudik et al., 2011), SnIPS, Self-Normalized Doubly Robust (SnDR) (Swaminathan & Joachims, 2015b), More Robust Doubly Robust Estimator (MRDR) (Farajtabar et al., 2018), and Distributionally Robust Policy Evaluation (DRPE) (Si et al., 2020). We include a variant of our proposed methods DM (R), shown in Equation 23, as a baseline by setting the density ratio in the robust regression to 1. More details about the baselines, including SnIPS, SnDR, IPS-GCS, and DM(R), can be found in Appendix B. Additionally, we provide the further comparison with other OPE methods in Appendix E.4, including SWITCH (Wang et al., 2017) and Doubly robust estimator with Shrinkage (DRoS) (Su et al., 2020), demonstrating the advantage of our method when integrated with various kinds of OPE approaches.